• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Eugene

    Log entry time 16:07:35 on October 8,2002

    Entry number 87634

    keyword=Moller measurement

    We started Moller measurement at 9:30.
    In an hour we discovered a problem with the DAQ. We made
    a control access at 11:00 to 12:00 to fix the problem. It turned out that
    the trigger module (TIR) was messing up the input ECL signals,
    a week ago this was the helicity and 30Hz syncro-signal, this
    time it was the ADC trigger and some others.
    Pluggin/unplugging the cables fixed the problem, but no clear reason
    has been found. One possible reason could be an extra signal
    plugged a week ago into this module. This signal is now moved
    to a different NIM-ELC converter.